# A Resolution Independent Neural Operator

### v2

@@ 3.1.2 Example 2: Nonlinear 1D Darcy’s Equation
The worst-case RINO predictions are shown in Figure 7(a), demonstrating visually acceptable accuracy for both the output function prediction and the input function reconstruction. Similarly the test sample at the 25th percentile of the error distribution is shown in Figure 7(b).
Figure 7: 1D Nonlinear Darcy’s Equation: RINO Prediction for two queried input functions at inference and for the testing dataset: (a) Worst-case sample with the highest error in output function prediction, (b) Sample at the 25th percentile of the error distribution (i.e., 75% of cases have less error than this value). The first column shows the queried input function data as ‘o’ and the masked points (not provided to the dictionary) as ‘x’, along with the reconstruction from dictionary learning. The second column shows the projection coefficients of the queried input function data onto the learned dictionary of basis function from Figure 4(c). The last column shows the output function prediction along with the true solution.
@@ 3.1.3 Example 3: Nonlinear 2D Darcy’s Equation
The RINO predictions corresponding to the worst-case test error for the output field is shown in Figure 8, demonstrating satisfactory error.
Figure 8: 2D Nonlinear Darcy’s Equation: RINO prediction for the worst-case output function $s(\boldsymbol{y})$ in the test dataset. The top left figure shows the queried input function data $u(\boldsymbol{x})$, with black circle symbols indicating the queried data. The middle top figure shows the input field reconstruction via the learned dictionary. The top right figure shows the error between the reconstructed input field and the ground truth. The bottom figures show the output field $s(\boldsymbol{y})$, the RINO prediction, and its corresponding errors.
3.1.4 Example 4: Burgers’ Equation
In this numerical example, we examined an initial boundary value problem in which the input and output functions of the operator have different domains and dimensions. Specifically, we looked at the parametric Burgers’ equation, as introduced in [42], with the random field initial condition $s(x)|_{t=0}$ as the input function and the space-time solution $s(x,t)$ as the output function:
$\frac{\partial s}{\partial t}+\frac{\partial s}{\partial x}=\nu\frac{\partial^{2}s}{\partial x^{2}};\quad(x,t)\in[0,1]^{2},$ (22)
@@ 3.1.4 Example 4: Burgers’ Equation
where periodic boundary conditions are assumed, and the periodic initial condition is a Gaussian random process with parameters similar to [42], and $\nu=0.01$. A spectral solver is used to obtain the solution field on a $101\times 100$ grid in the space-time domain for each input function, which is discretized on 101 uniformly spaced points. Following the procedure described earlier, we set $M_{\text{min}}=40$ and $M_{\text{max}}=70$ to generate random point clouds for the training and testing cases. The sizes of the training and test datasets are 1500 and 500, respectively. Notice that the input function samples are randomly discretized; however, the predictions are on the full solution grid for the results in this section, similar to other examples.
To showcase the flexibility of the RINO framework, we use predefined trunk basis vectors as the first 70 POD bases obtained from the solution output snapshots of the training dataset. It has been shown previously that using predefined optimal bases can facilitate the training of DeepONet and achieve faster convergence [23]. Two samples from the testing set are shown in Figure 9.
Figure 9: 2D Burger’s Equation: RINO prediction for two queried initial condition fields at inference and for the testing dataset: (a) Sample with the highest error in output function prediction $s(\boldsymbol{y})$, and (b) Sample at the 25th percentile of the error distribution (i.e., 75% of cases have less error than this value). The first column shows the ground truth solution. The second column shows the RINO prediction with the queried input initial condition points shown along the x-axis. The last column shows the point-wise error. The output domain $\boldsymbol{y}=(x,t)$ has spatio-temporal coordinates, where the horizontal axis represents space and the vertical axis represents time.
@@ 3.1.5 Discussion
After training each of the models listed above on the defined randomly sub-sampled data, we can perform inference during testing using any combination of input function sensors and at any desired points, thanks to the fully continuous nature of the RINO for processing both input and output functions. To examine the influence of discretization, we perform additional tests on regularly subsampled sensors for the input function at inference. A summary of our results is provided in Table 1. Rows in this table with a numeric value for the number of sensors (e.g., $M=100$) indicate performance during inference for input functions queried at structured and uniformly spaced sensor locations. Rows in the table denoted ‘random,’ correspond to test sets on the randomly sub-sampled discretizations described in Sections 3.1.1 – 3.1.4. The RINO used for predictions is trained using a data set of randomly sub-sampled points as described for each problem in Sections 3.1.1 – 3.1.4. Therefore, training error is reported only for the ‘random’ case.
@@
As the results suggest, the output function prediction is nearly insensitive to the discretization resolution of the queried input function for all PDE cases. However, in the 1D Nonlinear Darcy’s and Burgers’ equations, we observe a degradation in accuracy after some level of resolution coarsening (see the last row for each case). This occurs because, as anticipated, when the resolution falls below a certain threshold—particularly when it is significantly less than the problem’s characteristic length scale—the predictions become erroneous and thus resolution-dependent. In other words, these correspond to cases where a ReNO cannot be accurately identified due to under-sampling.
@@
Table 1: Resolution dependence study for each verification problem. RINO output function prediction errors are reported for queried input functions at various resolutions denoted by the Number of Sensors ($M$). Cases where a numerical value is prescribed correspond to fixed regular discretizations of the input. Cases denoted ‘random’ indicate a random subsampling of the input discretization as defined in Sections 3.1.1 – 3.1.4. Mean training/test errors and standard deviations are computed from five independent trials.
3.2 Demonstration Example: Nonlinear 2D Darcy’s Equation on Unstructured Meshes
@@
Operator Learning with unknown trunk
@@
Similar to the previous examples, using Algorithm 1, we learn 40 INR basis functions to effectively project an arbitrary input function point cloud onto the dictionary subspace with a satisfactory reconstruction error (see Appendix D). We then use this 40-dimensional latent representation for operator learning where the trunk and branch neural networks are unknown and are trained. The worst-case output function prediction and the case corresponding to the 25th percentile of the error distribution from the unseen test data are shown in  Figures 10 and 11, respectively. The results indicate a satisfactory training process and prediction capability. The full error distributions for the training and test data are provided in Appendix D.
@@
Figure 10: 2D Nonlinear Darcy’s Equation (Unstructured Mesh): RINO prediction with the worst-case output function, $s(\boldsymbol{y})$, in the test dataset. The top left figure shows the queried input function data $u(\boldsymbol{x})$. The middle top figure shows the input field reconstruction via learned dictionary. The top right figure shows the error between the reconstructed input field and the ground truth. The bottom figures show the true output field $s(\boldsymbol{y})$, the RINO prediction, and its corresponding errors.
@@
Figure 11: 2D Nonlinear Darcy’s Equation (Unstructured Mesh): RINO results for the test sample with an output function prediction error in the 25th percentile of the error distribution. The top left figure shows the queried input function data $u(\boldsymbol{x})$. The middle top figure shows the input field reconstruction via learned dictionary. The top right figure shows the error between the reconstructed input field and the ground truth. The bottom figures show the true output field $s(\boldsymbol{y})$, the RINO prediction, and its corresponding errors.
Operator Learning with predefined INR trunk
@@
As discussed in Section 2.1.1, the proposed dictionary learning algorithms can be applied in a similar manner for learning appropriate basis functions for the output function data. Once the dictionary is learned (offline), it can be utilized as a predefined trunk. Under such a setup, we only need to train the branch network. In this problem, we utilize Algorithm 2 to learn a dictionary of 54 INR basis functions that can sufficiently approximate the output function data. The worst-case output function prediction and the case corresponding to the 25th percentile of the error distribution from the unseen test data are shown in  Figures 12 and 13, respectively. The results indicate a satisfactory training process and prediction capability. The full error distributions for the training and test data are provided in Appendix D.
@@
Figure 12: 2D Nonlinear Darcy’s Equation (Unstructured Mesh) with Predefined INR Trunk Net: RINO prediction with the worst-case output function, $s(\boldsymbol{y})$, in the test dataset. The top left figure shows the queried input function data $u(\boldsymbol{x})$. The middle top figure shows the input field reconstruction via learned dictionary. The top right figure shows the error between the reconstructed input field and the ground truth. The bottom figures show the true output field $s(\boldsymbol{y})$, the RINO prediction, and its corresponding errors.
@@
Figure 13: 2D Nonlinear Darcy’s Equation (Unstructured Mesh) with Predefined Trunk Net: RINO results for the test sample with an output function prediction error in the 25th percentile of the error distribution. The top left figure shows the queried input function data $u(\boldsymbol{x})$. The middle top figure shows the input field reconstruction via learned dictionary. The top right figure shows the error between the reconstructed input field and the ground truth. The bottom figures show the true output field $s(\boldsymbol{y})$, the RINO prediction, and its corresponding errors.
@@ 4 Conclusion
In this work, we propose a resolution-independent neural operator (RINO) that modifies the DeepONet formulation with minimal architectural changes to directly work on (sufficiently rich) point clouds sampled from input and output functions without any underlying structural assumptions on the data structure. Specifically, we introduce a dictionary learning algorithm to identify appropriate continuous and fully differentiable basis functions for input function realizations. This allows us to construct an appropriate coordinate system (frame) to project input function point clouds onto, and use the resulting coordinates as the input function representation, which can be subsequently used in the vanilla DeepONet. The application of the introduced algorithm extends beyond the operator learning task and can be used for reconstruction tasks from incomplete or missing information, similar to GPOD. This algorithm can be considered a reconstruction method over function spaces rather than vector spaces, as is the case with GPOD or PCA.
@@
The introduced algorithm for finding a resolution-independent representation of the input function data has two additional advantages that “come for free”. First, if there is an underlying lower-dimensional structure for the realizations, the algorithm can identify it and reduce the dimensionality of the input function data. This was demonstrated in all our numerical examples. For instance, in one example, the data dimensionality was 100, and we achieved a relatively low reconstruction error using only around 10 basis functions. This compact data representation can directly reduce the size of the branch network. Moreover, such reduced dimensions are more suited for uncertainty quantification and propagation. Second, the reconstruction error can naturally serve as a good estimate of out-of-sample or distribution indicators, which can be leveraged for data acquisition and experiment design.

### v3

@@ 3.1.2 Example 2: Nonlinear 1D Darcy’s Equation
The worst-case RI-DeepONet predictions are shown in Figure 7(a), demonstrating visually acceptable accuracy for both the output function prediction and the input function reconstruction. Similarly the test sample at the 25th percentile of the error distribution is shown in Figure 7(b).
Figure 7: 1D Nonlinear Darcy’s Equation: RI-DeepONet Prediction for two queried input functions at inference and for the testing dataset: (a) Worst-case sample with the highest error in output function prediction, (b) Sample at the 25th percentile of the error distribution (i.e., 75% of cases have less error than this value). The first column shows the queried input function data as ‘o’ along with the reconstruction from dictionary learning. The second column shows the projection coefficients of the queried input function data onto the learned dictionary of basis function from Figure 4(c). The last column shows the output function prediction along with the true solution.
@@ 3.1.3 Example 3: Nonlinear 2D Darcy’s Equation
The RI-DeepONet predictions corresponding to the worst-case test error for the output field is shown in Figure 8, demonstrating satisfactory error.
Figure 8: 2D Nonlinear Darcy’s Equation: RI-DeepONet prediction for the worst-case output function $s(\boldsymbol{y})$ in the test dataset. The top left figure shows the queried input function data $u(\boldsymbol{x})$, with black circle symbols indicating the queried data. The middle top figure shows the input field reconstruction via the learned dictionary. The top right figure shows the error between the reconstructed input field and the ground truth. The bottom figures show the output field $s(\boldsymbol{y})$, the RI-DeepONet prediction, and its corresponding errors.
3.1.4 Example 4: Nonlinear Burgers’ Equation
In this numerical example, we examined an initial boundary value problem in which the input and output functions of the operator have different domains and dimensions. Specifically, we looked at the parametric nonlinear Burgers’ equation, as introduced in [60], with the random field initial condition $s(x)|_{t=0}$ as the input function and the space-time solution $s(x,t)$ as the output function:
$\frac{\partial s}{\partial t}+s\frac{\partial s}{\partial x}=\nu\frac{\partial^{2}s}{\partial x^{2}};\quad(x,t)\in[0,1]^{2},$ (22)
@@ 3.1.4 Example 4: Nonlinear Burgers’ Equation
where periodic boundary conditions are assumed, and the periodic initial condition is a Gaussian random process with parameters similar to [60], and $\nu=0.01$. A spectral solver is used to obtain the solution field on a $101\times 100$ grid in the space-time domain for each input function, which is discretized on 101 uniformly spaced points. Following the procedure described earlier, we set $M_{\text{min}}=40$ and $M_{\text{max}}=70$ to generate random point clouds for the training and testing cases. The sizes of the training and test datasets are 1500 and 500, respectively. Note that the input function samples are randomly discretized; however, the predictions are on the full solution grid for the results in this section, similar to other examples.
To showcase the flexibility of the RI-DeepONet framework, we use predefined trunk basis vectors as the first 70 POD bases obtained from the solution output snapshots of the training dataset (making it effectively RI-POD-DeepONet). It has been shown previously that using predefined optimal bases can facilitate the training of DeepONet and achieve faster convergence [35]. Two samples from the testing set are shown in Figure 9.
Figure 9: 2D Burger’s Equation: RI-DeepONet prediction for two queried initial condition fields at inference and for the testing dataset: (a) Sample with the highest error in output function prediction $s(\boldsymbol{y})$, and (b) Sample at the 25th percentile of the error distribution (i.e., 75% of cases have less error than this value). The first column shows the ground truth solution. The second column shows the RI-DeepONet prediction with the queried input initial condition points shown along the x-axis. The last column shows the point-wise error. The output domain $\boldsymbol{y}=(x,t)$ has spatio-temporal coordinates, where the horizontal axis represents space and the vertical axis represents time.
@@ 3.1.5 Discussion
After training each of the models listed above on the defined randomly sub-sampled data, we can perform inference during testing using any combination of input function sensors and at any desired points, thanks to the fully continuous nature of the RI-DeepONet for processing both input and output functions. To examine the influence of discretization, we perform additional tests on regularly subsampled sensors for the input function at inference. A summary of the results is provided in Table 1. Rows in this table with a numeric value for the number of sensors (e.g., $M=100$) indicate performance during inference for input functions queried at structured and uniformly spaced sensor locations. Rows in the table denoted ‘random,’ correspond to test sets on the randomly sub-sampled discretizations described in Sections 3.1.1 – 3.1.4. The RI-DeepONet used for predictions is trained using a data set of randomly sub-sampled points as described for each problem in Sections 3.1.1 – 3.1.4. Therefore, training error is reported only for the ‘random’ case.
@@
As the results suggest, the output function prediction is nearly insensitive to the discretization resolution of the queried input function for all PDE cases. However, in the 1D nonlinear Darcy’s and nonlinear Burgers’ equations, we observe a degradation in accuracy after some level of resolution coarsening (see the last row for each case). This occurs because, as anticipated, when the resolution falls below a certain threshold—particularly when it is significantly less than the problem’s characteristic length scale—the predictions become erroneous and thus resolution-dependent. In other words, these correspond to cases where a ReNO cannot be accurately identified due to under-sampling.
@@
Table 1: Resolution dependence study for each verification problem. RI-DeepONet output function prediction errors are reported for queried input functions at various resolutions denoted by the Number of Sensors ($M$). Cases where a numerical value is prescribed correspond to fixed regular discretizations of the input. Cases denoted ‘random’ indicate a random subsampling of the input discretization as defined in Sections 3.1.1 – 3.1.4. Mean training/test errors and standard deviations are computed from five independent trials.
3.2 Demonstration Example 1: Nonlinear 2D Darcy’s Equation on Unstructured Meshes
@@
Operator Learning with RI-DeepONet.
@@
Similar to the previous examples, using Algorithm 1, we learn 40 INR basis functions to effectively project an arbitrary input function point cloud onto the dictionary subspace with a satisfactory reconstruction error (see Appendix D). We then use this 40-dimensional latent representation for operator learning where the trunk and branch neural networks are unknown a priori. The evolution of the loss function during the operator learning stage is depicted in Figure 10(a). The distribution of output function prediction errors after training is plotted in Figure 10(b).
Figure 10: 2D Darcy’s Equation on Unstructured Meshes with RI-DeepONet: (a) Loss function during optimization iterations for operator learning. (b) Distribution of output function prediction errors for the training and testing datasets after training.
The worst-case output function prediction and the case corresponding to the 25th percentile of the error distribution from the unseen test data are shown in  Figures 11 and 12, respectively. The results indicate a satisfactory training process and prediction capability.
@@
Figure 11: 2D Nonlinear Darcy’s Equation (Unstructured Mesh): RI-DeepONet prediction with the worst-case output function, $s(\boldsymbol{y})$, in the test dataset. The top left figure shows the queried input function data $u(\boldsymbol{x})$. The middle top figure shows the input field reconstruction via learned dictionary. The top right figure shows the error between the reconstructed input field and the ground truth. The bottom figures show the true output field $s(\boldsymbol{y})$, the RI-DeepONet prediction, and its corresponding errors.
@@
Figure 12: 2D Nonlinear Darcy’s Equation (Unstructured Mesh): RI-DeepONet results for the test sample with an output function prediction error in the 25th percentile of the error distribution. The top left figure shows the queried input function data $u(\boldsymbol{x})$. The middle top figure shows the input field reconstruction via learned dictionary. The top right figure shows the error between the reconstructed input field and the ground truth. The bottom figures show the true output field $s(\boldsymbol{y})$, the RI-DeepONet prediction, and its corresponding errors.
Operator Learning with
@@
As discussed in Section 2.1.1, the proposed dictionary learning algorithms can be applied in a similar manner for learning appropriate basis functions for the output function data. Once the dictionary is learned (offline), it can be utilized as a predefined basis for the output. Under such a setup, we only need to train a single neural network between embedding spaces as a neural operator. In this problem, we utilize Algorithm 2 to learn a dictionary of 37 INR basis functions that can sufficiently approximate the output function data. The full error distributions of the learned output basis functions for the training and test data are provided in Appendix D.
The evolution of the loss function during the operator learning stage for the RINO case where we utilized the learned output basis dictionary is depicted in Figure 13(a). The distribution of output function prediction errors after training is plotted in Figure 13(b).
Figure 13: 2D Darcy’s Equation on Unstructured Meshes with RINO: (a) Loss function during optimization iterations for operator learning. (b) Distribution of output function prediction errors for the training and testing datasets after training.
The worst-case output function prediction and the case corresponding to the 25th percentile of the error distribution from the unseen test data are shown in  Figures 14 and 15, respectively. A successful training with satisfactory accuracy is confirmed by the results for the RINO case where we indirectly learned a mapping between the coefficients of two sets of weakly orthogonal basis functions: one corresponding to the input function data and the other to the output function data, which were identified separately and offline.
@@
Figure 14: 2D Nonlinear Darcy’s Equation (Unstructured Mesh) with Predefined INR Output Basis: RINO prediction with the worst-case output function, $s(\boldsymbol{y})$, in the test dataset. The top left figure shows the queried input function data $u(\boldsymbol{x})$. The middle top figure shows the input field reconstruction via learned dictionary. The top right figure shows the error between the reconstructed input field and the ground truth. The bottom figures show the true output field $s(\boldsymbol{y})$, the RINO prediction, and its corresponding errors.
@@
Figure 15: 2D Nonlinear Darcy’s Equation (Unstructured Mesh) with Predefined INR Output Basis: RINO results for the test sample with an output function prediction error in the 25th percentile of the error distribution. The top left figure shows the queried input function data $u(\boldsymbol{x})$. The middle top figure shows the input field reconstruction via learned dictionary. The top right figure shows the error between the reconstructed input field and the ground truth. The bottom figures show the true output field $s(\boldsymbol{y})$, the RINO prediction, and its corresponding errors.
From these results and the error plots presented for the case RI-DeepONet, we see that the RI-DeepONet and RINO results have comparable accuracy. However, the RINO has the distinct advantage of learning the operator in the compact embedding space such that the input and output bases do not have redundancies (due to their orthogonality). The result of this is that the neural operator itself is much more compact – having far fewer parameters. Where the RI-DeepONet uses MLP with layers [40, 50, 100], the RINO uses MLP with layers [40, 40, 37] to achieve comparable accuracy (see Appendix E). Other potential advantages of RINO over RI-DeepONet, particularly in terms of improved generalization capability, are also discussed in Appendix G.
3.3 Demonstration Example 2: Solid Mechanics
In the absence of body forces, the balance of linear momentum, combined with elastic constitutive laws, can be expressed as follows:
$\frac{\partial s_{ij}}{\partial x_{i}}$ $=0,\quad s_{ij}=C_{ijkl}(\boldsymbol{x})\epsilon_{kl},\quad C_{ijkl}(\boldsymbol{x})=\lambda(\boldsymbol{x})\delta_{ij}\delta_{kl}+\mu(\boldsymbol{x})\left(\delta_{ik}\delta_{jl}+\delta_{il}\delta_{jk}\right),$ (24)
where $s_{ij}$ denotes symmetric stress tensor components, $\epsilon_{ij}$ denotes infinitesimal strain tensor components, $\mathbb{C}$ denotes elasticity tensor and Lame’s first parameter $\lambda$ and shear modulus $\mu$ are related to Elastic modulus and Poisson’s ratio. Poisson’s ratio is set to a constant value of 0.3 throughout the domain. However, the elastic modulus modeled as a Gaussian random field with mean 15MPa, correlation length scale 2.5mm and standard deviation 0.25MPa, which varies spatially across the domain and differs among realizations. We consider a square plate with a side length of 1cm subjected to uniaxial loading along the vertical axis under plane stress conditions. The applied traction is 0.42MPa. We generate data on structured quadrilateral finite element meshes of varying densities, where the number of elements along each side of the square plate is a uniformly distributed random variable between 40 and 60. For illustration, three realizations are plotted in Figure 16.
Figure 16: Solid Mechanics Model: Finite element nodal values of the elastic modulus field for three realizations, with varying mesh densities across the realizations. Units are in MPa.
In this example, we aim to learn the stress fields corresponding to a given elastic modulus field. The proposed method is applied separately to each component of the stress tensor, namely $s_{11}$, $s_{12}$, and $s_{22}$. We apply the RINO method described in Section 2.1.1, where two separate sets of basis functions are learned: one for the input function and another for the output function. The basis functions are learned using the procedure outlined in Algorithm 2 with network architecture is described in Appendix E.
As described earlier, once an appropriate set of basis functions for the input and output fields is identified, the operator learning task reduces to mapping the input embeddings (input basis coefficients) to the output embeddings (output basis coefficients). This mapping is parameterized using a MLP (see Appendix E).
The worst-case relative error for stress component predictions on unseen test cases is plotted in Figure 17, demonstrating good accuracy. To further validate this, the relative mean square error across all training and test realizations is shown in Figure 18. The results suggest that not only is high accuracy achieved, but good generalization is also evident.
Figure 17: Solid Mechanics Model: Worst case stress component predictions among test realizations: (a) $\sigma_{11}$, (b) $\sigma_{12}$, (c) $\sigma_{22}$.
Figure 18: Relative error distribution of stress component predictions among 4000 training and 1000 test realizations: (a) $\sigma_{11}$, (b) $\sigma_{12}$, (c) $\sigma_{22}$.
@@ 4 Conclusion
In this work, we propose a resolution-independent neural operator (RINO) that builds from the DeepONet formulation to directly work on (sufficiently rich) point clouds sampled from input and output functions without any underlying assumptions on the data structure. Specifically, we introduce a dictionary learning algorithm to identify appropriate continuous and fully differentiable basis functions for input and output function realizations. In the context of DeepONet, this allows us to construct an appropriate coordinate system (frame) to project input function point clouds onto, and use the resulting coordinates as the input function representation in the vanilla DeepONet; thus creating a resolution-independent DeepONet (RI-DeepONet). In fact, by applying the proposed dictionary learning algorithm separately to the input and output functions, we demonstrate that the operator learning task reduces to learning a mapping function between the coefficients of the input and output basis functions. This can be viewed as a new operator learning framework termed the RINO. Moreover, we show that for certain problems, the approximate operator in the constructed latent spaces significantly simplifies the operator learning optimization process – e.g., a simple linear transformation between the embeddings. The application of the proposed algorithm extends beyond the operator learning task and can be used for reconstruction tasks from incomplete or missing information, similar to GPOD. This algorithm can be considered a reconstruction method over function spaces rather than vector spaces, as is the case with GPOD or PCA.
@@
The introduced algorithm for finding a resolution-independent representation of the input function data has two additional advantages that “come for free”. First, if there is an underlying lower-dimensional structure for the realizations, the algorithm can identify it and reduce the dimensionality of the input/output function data. This was demonstrated in all of the numerical examples. For instance, in one example, the data dimensionality was 100, and we achieved a relatively low reconstruction error using only 10 basis functions. This compact data representation can directly reduce the size of the operator network. Moreover, such reduced dimensions are more suited for uncertainty quantification and propagation. Second, the reconstruction error can naturally serve as a good estimate of out-of-sample or distribution indicators, which can be leveraged for data acquisition and experiment design.